# Post-Incorporating Code Structural Knowledge into Pretrained Models via ICL for Code Translation

### v1

@@ IV-D Empirical Results and Analysis
[…]
question-circle RQ8: What is the impact of our method on failure frequencies?
@@
[…]
V Threat Analysis
@@
VI Related Works
VI-A Code translation
@@
Automated code translation aids in the migration of codebases across programming languages, significantly enhancing productivity in software development. Early rule-based translation methods have been extensively explored [41, 42, 43, 44]. Furthermore, phrase-based statistical machine translation techniques have shown notable advances [45, 46, 47, 48, 49, 50]. Due to the success of deep learning, deep code translation methods have extended the SOTA further [51, 52, 53, 54, 55].
@@
Prior research has shown that, unlike the grammatical structure of natural language, the syntactic structure of source code is stricter and more intricate, directly reflecting the program’s behavior [6, 56, 57, 58, 9]. Previous studies have demonstrated that explicitly incorporating structural knowledge into models enhances their perceptual understanding of code, thereby aiding software mining tasks [11, 12], including code translation [10, 6]. For example, modeling structural knowledge from AST [5, 6, 7], control flow graphs (CFG) [8, 9], and data flow graphs (DFG) [10]. However, these syntax-aware methods depend on meticulously crafted model architectures and loss functions tailored through training.
@@
Recently, LLMs, scaled to hundreds of billions of parameters, have exhibited remarkable performance across various natural language processing tasks [59, 60, 61]. By training on vast code corpora, several LLMs have achieved impressive performance, such as GPT series [62, 38], LLaMA series [37, 36], Qwen series [32, 33], and Gemma series [63, 34]. Furthermore, to specifically advance in code tasks, some LLMs are pre-trained and continuously trained with more code data and code-specific tasks, including CodeLLaMA [3], CodeGemma [64], Qwen2.5-Coder [33], Starcoder series [4, 65], and DeepSeek-Coder series [1, 35]. However, mainstream LLMs often overlook code structural knowledge, whose effectiveness has been validated in prior methods. Directly combining the carefully designed model architectures and loss functions of classic syntax-aware methods with LLMs is both time- and resource-intensive, even be prohibited for proprietary models. Therefore, research on how to post-incorporate code structural knowledge into off-the-shelf LLMs is limited but crucial.
VI-B In-context Learning
@@
Recent studies have shown that LLMs exhibit strong ICL abilities as model and data sizes scale up [66, 67, 68]. The essence of ICL lies in learning by analogy, requiring a prompt context composed of a few demonstration exemplars. By concatenating these exemplars with a query, LLMs can make better predictions without updating parameters, differentiating ICL from traditional model tuning. Recently, numerous studies have attempted to leverage ICL for complex tasks, including natural language understanding [14], reasoning [69, 70, 71], text generation [23, 72, 22]. Additionally, several studies aim to provide theoretical justifications and insights into how LLMs learn from a few in-context demonstration exemplars [73, 74]. Despite significant progress, ICL still faces challenges and requires further exploration, such as its sensitivity to the selection of demonstration exemplars [75].
@@
Various objectives for ICL exemplar selection and tailoring have been explored [76, 77, 16, 13]. Mostly ICL exemplar selection typically relies on point-wise similarity [14, 15, 16], where a matching score is computed between each candidate exemplar and the input query, followed by the independent selection of the top-$k$ exemplars to construct the context [13]. Although these methods are straightforward and fast, they fail to consider compositional relationships among the selected $k$ exemplars, often leading to redundancy caused by homogeneity. Some studies aim to ensure diversity by clustering exemplars into $k$ groups, then selecting the most similar exemplars from each group for a given query [17, 18, 19]. Besides, some works [20, 21] introduce an auxiliary loss to regularize diversity. However, additional diversity loss may hinder the optimization of relevance, and accounting for compositional relationships in diversity remains challenging. Some methods attempt to combine weighted similarity and diversity, but adaptively balancing these two objectives is difficult. Moreover, these methods do not focus on code translation and are typically syntax-unaware, neglecting the critical syntactic information inherent in source code. Additionally, [21] attempts to account for local structure coverage in ICL. However, it only considers the structure between a parent node and its child nodes, and the algorithm is rough, needing to work together with complex similarity and diversity regularization losses. In contrast, our method is grounded in information theory, providing greater precision and generality. We consider a more comprehensive subtree structure, and our algorithm can be applied directly, guaranteed by submodular theory, without requiring additional losses.
VII Conclusion and Future Work
@@
Code translation aids code migration between languages, boosting productivity in software development. This work focuses on post-incorporating code structural knowledge into pre-trained models via ICL for code translation. We revisit exemplar selection in ICL from an information-theoretic perspective, showing that maximizing information coverage provides a more precise and general solution than traditional methods. We introduce CAST, a surrogate measure based on the maximum subtree coverage ratio of AST. We show that NP-hard CAST maximization is a submodular maximization problem and propose a greedy algorithm with a $(1-1/e)$-approximate solution of polynomial time complexity. Our method is the first training-free, model-agnostic approach to integrate code structural knowledge into existing LLMs at test time, and both theory and experiments prove their effectiveness and efficiency. This work provides two key insights: 1) Code structural knowledge can be post-incorporated into pre-trained LLMs during inference despite training neglect. 2) Scaling model size or data doesn’t lead to code structural knowledge, highlighting the need to consider code syntax in LLMs.
@@ Acknowledgments
This research is supported by the National Natural Science Foundation of China (NSFC) under grant numbers 62076121 and 61921006, the Major Program (JD) of Hubei Province (2023BAA024), and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX24_0301).

### v2

@@ IV-D Empirical Results and Analysis
ȷ Finding 7: Both CAST-F and CAST-A demonstrate strong robustness and stability across varying hyperparameter settings.
RQ8: What is the impact of our method on failure frequencies?
@@
ȷ Finding 8: CAST-F effectively reduces failure frequencies, achieving higher accuracy and robustness compared to baseline strategies, particularly in minimizing “Wrong Answer” outcomes and maximizing “Accept” rates.
TABLE VI: Results of Qwen2-7B for 5-shot Code summarization upon different ICL strategies.
Strategy Java Python PHP Go Javascript Ruby
BLEU Rouge-L BLEU Rouge-L BLEU Rouge-L BLEU Rouge-L BLEU Rouge-L BLEU Rouge-L
Random 2.25 9.01 5.29 14.80 12.30 10.78 4.60 14.93 2.33 5.45 5.02 13.13
BM25 2.27 9.60 5.92 14.74 17.65 16.55 4.48 14.24 2.86 6.03 5.49 12.50
LD 2.06 9.39 4.47 13.45 16.27 16.14 5.43 15.63 3.11 7.83 4.82 13.00
CodeBERT 2.43 9.56 5.92 15.07 17.41 15.05 4.34 13.35 2.80 5.87 5.99 13.14
Diversity 2.30 9.39 5.23 13.00 14.52 13.40 3.88 12.80 2.43 5.86 5.25 13.34
CAST-F 2.42 9.90 6.73 16.57 17.58 18.65 5.52 17.68 4.81 9.12 6.71 14.47
RQ9: Could CAST be extended to support programming languages and tasks beyond code translation?
To showcase the broader applicability of CAST beyond code translation, we conduct an additional experiment on code summarization. Code summarization generates natural language descriptions from source code snippets, a crucial task in software maintenance and comprehension. We use the widely used dataset provided by the CodeXGLUE [36] team for this task. We evaluate the performance of our method, using the widely adopted Qwen2-7B model across six programming languages: Java, Python, PHP, Go, JavaScript, and Ruby. The evaluation metrics include BLEU [45] and Rouge-L scores, assessing the quality of generated summaries against reference descriptions. Rouge-L computes an F-score based on the longest common subsequence, comparing the similarity between two texts.
Tab. VI presents results comparing CAST-F with baseline exemplar selection strategies. CAST-F achieves the highest BLEU and Rouge-L scores across nearly all languages and metrics, demonstrating its superior ability to select exemplars that enhance the model’s summarization quality. This consistent improvement underscores that CAST’s structural coverage principle captures critical code characteristics essential for generating accurate and informative summaries. These findings emphasize CAST’s generality and robustness as an exemplar selection metric. By capturing structural coverage effectively, CAST facilitates improved task-specific knowledge transfer for LLMs in code-related tasks.
ȷ Finding 9: The experiment on code summarization broadens the scope of our approach beyond code translation, establishing CAST as a versatile tool for enhancing code retrieval and generation tasks in various software engineering scenarios.
RQ10: How do different database sizes and noisy data affect the method’s performance?
We evaluated the impact of database size variations and noise injection on CAST performance using the TransCoder dataset Java-Python task under Qwen-7B. To assess database size effects, we randomly sampled subsets at 75%, 50%, and 25% of the full database as ICL demonstration exemplars, simulating scenarios with limited available data. For noise injection, we implemented three perturbation types to simulate real-world code imperfections:
  • Syntax Error: Random deletion or replacement of punctuation marks (semicolons, parentheses, quotes), introducing minor cross-language syntax errors.
  • Identifier Rename: Variable name obfuscation or random replacement, breaking identifier consistency and simulating naming errors.
  • Statement Shuffle: Random reordering of consecutive statements within code snippets, mimicking structural disorder in semantically related blocks.
As shown in Fig. 8a, reducing the database size for exemplar sampling leads to gradual accuracy degradation. Performance decreases from 69.01% (full database) to 67.08% (25% sampling), indicating that while CAST maintains robustness under reduced exemplar availability, performance diminishes as database information decreases. Fig. 8b demonstrates performance degradation under different noise types. Accuracy drops from 69.01% (no noise) to 66.20% (statement shuffle), with syntax errors and identifier renaming showing intermediate impact (67.25% and 67.08% respectively). These results indicate CAST’s relative resilience to syntax and naming perturbations but higher sensitivity to structural disruptions.
ȷ Finding 10: CAST performance degrades with reduced database size (2.93% drop at 25% sampling) and structural noise (2.81% loss with statement shuffle). Adequate database scale and code quality are essential for optimal performance.
Figure 7: A case study (qualitative analysis) of code translation comparing similarity-only method (i.e., LD) and our CAST-F, based on Qwen2-7B.
(a) Database Size Impact
(b) Database Noise Impact
Figure 8: Impact of database characteristics on computation accuracy. Fig. 8a Database size effect with random sampling. Fig. 8b Performance under different noisy database.
V Case Study
To understand why traditional similarity-based retrieval fails to provide optimal exemplars for code translation, we conduct a case study examining how different retrieval strategies affect translation quality. The fundamental limitation of similarity-only approaches is their tendency to retrieve exemplars with high similarity while overlooking sufficient knowledge required for the current translation task. Fig. 7 shows this through a C++ to Python translation task about findpos function that processes string-based position calculations.
The similarity-only retrieval (LD) operates purely on lexical similarity, leading to the selection of highly correlated but redundant exemplars that cannot provide sufficient knowledge for the current translation task. The case shows that the translated code fails in string boundary checking, resulting in an IndexError: string index out of range, where correctly translating this construct requires knowledge of the difference between C++ null-terminator-based string handling (’\0’) and Python’s length-based indexing (len()). However, the retrieved exemplars fail to provide this critical cross-language knowledge for properly handling the string termination construct while (n[i] != ’\0’).
In contrast, CAST retrieval functions as a surrogate objective for information coverage across syntactic structures, actively seeking exemplars that provide the knowledge required for the current translation task. By retrieving exemplars that collectively cover diverse syntactic structural patterns, CAST enables the model to discover the “aha insight” that Python requires length-based indexing (while i < len(n)). This demonstrates how CAST’s information coverage principle guides the model toward semantically correct translations by ensuring comprehensive knowledge representation in the retrieved exemplar set. The success stems from CAST’s ability to prioritize knowledge completeness over superficial similarity, leading to more robust code translation.
VI Threat Analysis
@@
VII Related Works
VII-A Code translation
@@
Automated code translation aids in the migration of codebases across programming languages, significantly enhancing productivity in software development. Early rule-based translation methods have been extensively explored [47, 48, 49, 50]. Furthermore, phrase-based statistical machine translation techniques have shown notable advances [51, 52, 53, 54, 55, 56]. Due to the success of deep learning, deep code translation methods have extended the SOTA further [57, 58, 59, 60, 61, 62].
@@
Prior research has shown that, unlike the grammatical structure of natural language, the syntactic structure of source code is stricter and more intricate, directly reflecting the program’s behavior [9, 63, 64, 65, 12]. Previous studies have demonstrated that explicitly incorporating structural knowledge into models enhances their perceptual understanding of code, thereby aiding software mining tasks [14, 15], including code translation [13, 9]. For example, modeling structural knowledge from AST [8, 10, 66, 67], control flow graphs (CFG) [11, 12], and data flow graphs (DFG) [13]. However, these syntax-aware methods rely on meticulously crafted model architectures and loss functions that are tailored through training.
@@
Recently, LLMs, scaled to hundreds of billions of parameters, have exhibited remarkable performance across various natural language processing tasks [68, 69, 70], such as GPT [71, 43], LLaMA [41], Qwen [38], etc. The LLM-powered code translation has achieved significant progress in recent years [72, 34, 73, 74]. For example, Yang et al. [34] proposed UniTrans, which introduces auto-generated test cases throughout the whole code translation framework via translation augmentation and repair. Although UniTrans injects test information and achieves certain effectiveness, it overlooks the role of intrinsic code structure information in assisting the model to understand code semantics. In fact, mainstream methods based on LLMs often overlook code structural knowledge, whose effectiveness has been validated in prior studies [12, 65, 11]. Directly combining the carefully designed model architectures and loss functions of classic syntax-aware methods with LLMs is both time- and resource-intensive, and may even be prohibited for proprietary models. Therefore, research on how to post-incorporate code structural knowledge into off-the-shelf LLMs is limited but crucial.
VII-B In-context Learning
@@
Recent studies have shown that LLMs exhibit strong ICL abilities as model and data sizes scale up [75, 76, 77]. The essence of ICL lies in learning by analogy, requiring a prompt context composed of a few demonstration exemplars. By concatenating these exemplars with a query, LLMs can make better predictions without updating parameters, differentiating ICL from traditional model tuning. Recently, numerous studies have attempted to leverage ICL for complex tasks, including natural language understanding [17], reasoning [78, 79, 80], text generation [26, 81, 25]. Additionally, several studies aim to provide theoretical justifications and insights into how LLMs learn from a few in-context demonstration exemplars [82, 83]. Despite significant progress, ICL still faces challenges and requires further exploration. such as its sensitivity to the selection of demonstration exemplars [84].
@@
Various objectives for ICL exemplar selection and tailoring have been explored [85, 86, 19, 16]. Mostly ICL exemplar selection typically relies on point-wise similarity [17, 18, 19, 66, 67], where a matching score is computed between each candidate exemplar and the input query, followed by the independent selection of the top-$k$ exemplars to construct the context [16]. Although these methods are straightforward and fast, they fail to consider compositional relationships among the selected $k$ exemplars, often leading to redundancy caused by homogeneity. Some studies aim to ensure diversity by clustering exemplars into $k$ groups, then selecting the most similar exemplars from each group for a given query [20, 21, 22]. Besides, some works [23, 24] introduce an auxiliary loss to regularize diversity. However, additional diversity loss may hinder the optimization of relevance, and accounting for compositional relationships in diversity remains challenging. Some methods attempt to combine weighted similarity and diversity, but adaptively balancing these two objectives is difficult. Moreover, these methods do not focus on code translation and are typically syntax-unaware, neglecting the critical syntactic information inherent in source code. Additionally, [24] attempts to account for local structure coverage in ICL. However, it only considers the structure between a parent node and its child nodes, and the algorithm is rough, needing to work together with complex similarity and diversity regularization losses. In contrast, our method is grounded in information theory, providing greater precision and generality. We consider a more comprehensive subtree structure, and our algorithm can be applied directly, guaranteed by submodular theory, without requiring additional losses.
VIII Conclusion and Future Work
@@
Code translation aids code migration between languages, boosting productivity in software development. This work focuses on post-incorporating code structural knowledge into pre-trained models via ICL for code translation. We revisit exemplar selection in ICL from an information-theoretic perspective, showing that maximizing information coverage provides a more precise and general solution than traditional methods. We introduce CAST, a surrogate measure based on the maximum subtree coverage ratio of AST. We show that NP-hard CAST maximization is a submodular maximization problem and propose a greedy algorithm with a (11/e)(1-1/e)-approximate solution of polynomial time complexity. Our method is the first training-free, model-agnostic approach to integrate code structural knowledge into existing LLMs at test time, and both theory and experiments prove their effectiveness and efficiency. This work provides two key insights: 1) Code structural knowledge can be post-incorporated into pre-trained LLMs during inference despite training neglect. 2) Scaling model size or data doesn’t lead to code structural knowledge, highlighting the need to consider code syntax in LLMs.
@@ Acknowledgments
This research is supported by the National Natural Science Foundation of China (NSFC) under grant number 62076121, the Major Program (JD) of Hubei Province (2023BAA024).